Find the value of "k" if f(x) = 3x2 + 6x + 1 is put into general form.
a. -2
b. -1
c. 1
d. 3
e. 4

Answer by uma(370)

f(x) = 3x^2 + 6x + 1
Let y = 3x^2 + 6x + 1
= 3(x^2 + 2x) + 1
= 3(x^2 + 2x + 1 - 1) + 1 [completing the square]
= 3(x^2 + 2x + 1) -3 + 1
= 3(x+1)^2 - 2
Comparing this with the standard form, y = a(x-h)^2 + k we get,
k = -2

Good Luck!!!